OpenStudy (arbyj):

The length of a poster is 7 in greater than the width. The perimeter is 114 in. Find the width and the length. 4w+14=114 is the equation. I got 20 as the width. What's the length?

OpenStudy (ddcamp):

"The length of a poster is 7 in greater than the width"

OpenStudy (arbyj):

@DDCamp So I multiply 7 x 20?

OpenStudy (ddcamp):

Let's look back at the equation you found for the area: 4w+14=114 How did you get that?

OpenStudy (arbyj):

My teacher had this answer bank set up for us and I assumed it would be this equation

OpenStudy (ddcamp):

l = w+7 P = 2w + 2l = 2w + 2(w+7) = 2w + 2w + 14 = 4w + 14

OpenStudy (ddcamp):

From that equation, we find w: 4w+14 = 114 4w = 100 w = 25

OpenStudy (arbyj):

Oh dang, I did the math wrong, I see

OpenStudy (ddcamp):

Then we find l from the original problem. The first line says that the length is seven inches longer than the width, or: l = w+ 7 l = 25+7 l = 32

OpenStudy (arbyj):

Oh ok